OpenStudy (anonymous):

A wave is modeled with the function y = 1/2 sin 3 theta. Describe the graph of this function, including its period, amplitude, and points of intersection with the x-axis.

OpenStudy (anonymous):

\[y = \frac{ 1 }{ 2 } \sin 3 \theta \]

OpenStudy (anonymous):

if someone can help me figure out how to graph it, i can find the period, amplitude, and points of intersection with the x-axis myself.

OpenStudy (anonymous):

the amplitude is \(\frac{1}{2}\) because of the coefficient out front

OpenStudy (anonymous):

hope this is clear, since \(\sin(x)\) goes from \(-1\) to \(1\) this means \(\frac{1}{2}\sin(x)\) goes from \(-\frac{1}{2}\) to \(\frac{1}{2}\)

OpenStudy (anonymous):

as for the period, the period of \(\sin(bx)\) is \(\frac{2\pi}{b}\) in this case the period will therefore be \(\frac{2\pi}{3}\)

OpenStudy (anonymous):

you can solve for the \(x\) intercepts, by setting the various zeros of \(\sin(x)\) equal to \(3x\) for example \[3x=0\iff x=0\] \[3x=\pi\iff x=\frac{\pi}{3}\] \[3x=2\pi\iff x=\frac{2\pi}{3}\] and so on

OpenStudy (anonymous):

as for a graph, i would use this http://www.wolframalpha.com/input/?i=1%2F+2*sin%283x%29

OpenStudy (anonymous):

I found some of the answers myself (finally figured it out!) I had these: Amplitude: 1/2 Period: 2pi/3 Intersections: Infinitely many, but pi/3, 2pi/3, so on and so forth. Hopefully those were right.

OpenStudy (anonymous):

From what you said, they look okay.

OpenStudy (anonymous):

yes, you are correct

OpenStudy (anonymous):

Awesome! Thanks for your help.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@satellite73 i don't get it :C

OpenStudy (anonymous):

how you get b